Magentic mument of an ions of iron is 5.92BM calculate unpaired electrons and show electronic configration

Grade:12

2 Answers

Tejas Shripal
6 years ago
Magnetic moment = \sqrt{n(n+2)}
5.92 = \sqrt{n(n+2)}
Square on both sides,
35 = n(n+2)
n^{2}+2n -35=0
After solving this you get n=5 , -7.
And since n is an postive integer ,we take 5 into consideration. The number of unpaired electrons is 5.
 
Tejas Shripal
6 years ago
And about electronic config, Fe has an electronic config [Ar] 3d6 4s2, Here only 4 electrons are upaired. By the question 5 are unpaired. fe+3 has 5 unpaired e- and its electronic config is 1s2 2s2 2p6 3s2 3p6 3d5.